OpenStudy (anonymous):

Use the matrix tool to sovle the system of equation. Enter the answer as an ordered pair. 6x+4y=52 4x+7y=52

OpenStudy (anonymous):

How can I use the matrix tool?

OpenStudy (anonymous):

[6 4 52] [4 7 52] 6*R2-4*R1----R2 [6 4 52] [0 26 104] y=4 x=6 (6,4)

OpenStudy (anonymous):

[6 4 52] row 1 divide by 6 [1 0.67 8.67] [4 7 52] row 2 divide by 4 [ 1 1.75 13 ] [1 1.75 13] subtract this to row2 [1 1.75 13] [1 .67 8.67] to get [0 1.08 4.33] divide R2 by 1.08 [ 1 1.75 13] [ 0 1 4 ] here y=4 then by back sub x+ 1.75y =13 x+ 1.75(4)=13 x=6 therefore x=6,y=4 ans